Let (a,b,c ) not equal to 0 be real constants. Show that the ellipsoid
E =((x,y,z) ∈R^3 : x^2/a^2 + y^2/b^2 + z^2/c^2 = 1) is a smooth surface.
1
Expert's answer
2020-10-14T13:55:45-0400

We can take "a,b,c>0."

We cover the ellipsoid by smooth charts of the form

"(x,y)\mapsto (x,y,c\sqrt{1-\frac{x^2}{a^2 }-\frac{y^2}{b^2 }})" and "(x,y)\mapsto (x,y,-c\sqrt{1-\frac{x^2}{a^2 }-\frac{y^2}{b^2 }})"

where "\{(x,y)\in(-a,a)\times(-b,b)\}"

Similarly, "(y,z)\mapsto (a\sqrt{1-\frac{z^2}{c^2 }-\frac{y^2}{b^2 }},y,z)" and "(y,z)\mapsto (-a\sqrt{1-\frac{z^2}{c^2 }-\frac{y^2}{b^2 }},y,z)"and lastly"(x,z)\mapsto (x,b\sqrt{1-\frac{z^2}{c^2 }-\frac{x^2}{a^2 }},z)" and "(x,z)\mapsto (x,b\sqrt{1-\frac{z^2}{c^2 }-\frac{x^2}{a^2 }},z)"These charts are smooth clearly, also these maps takes into account of all possible values of the ellipsoid since on the ellipsoid at least one of "x,y \ or z \neq 0." Hence is covered in the first two "(z\neq 0)" second two "(x\neq 0)" or last two "(y\neq 0)" .

Hence the ellipsoid can be covered by smooth atlas and hence is a smooth surface.